OpenStudy (xapproachesinfinity):

how to get this sum: \[\sum_{1}^{\infty}(-1)^{n+1}\frac{1}{n^5}\] I know it converges by the alternating series test but can't find a way to for the sum

myininaya (myininaya):

hmm wolfram mentioned something about the riemann zeta function you know anything about that function?

OpenStudy (xapproachesinfinity):

yeah i checked the wolfram, but we didn't do any do at all

OpenStudy (xapproachesinfinity):

of that *

OpenStudy (xapproachesinfinity):

reimann zeta of 5 is computed with integral yes

geerky42 (geerky42):

\[\sum_{n=1}^{\infty}(-1)^{n+1}\frac{1}{n^5}=\left(\sum_{n=1}^{\infty}\frac{1}{(2n-1)^5}\right)-\left(\sum_{n=1}^{\infty}\frac{1}{(2n)^5}\right)\]Yes?

OpenStudy (xapproachesinfinity):

hmm didn't get where did you got that lol

geerky42 (geerky42):

For every odd n, \((-1)^{n+1} = 1\) and for every even n, \((-1)^{n+1} = -1\), right? So what I did was arrange these terms so that positive and negative terms are grouped separately.

myininaya (myininaya):

After you said I was thinking telescoping series but that looks pretty ugly to me :(

myininaya (myininaya):

is that what you were going for ?

geerky42 (geerky42):

Hmm I have no idea what to do next XD I was basically suggested to arrange terms to perhaps make this problem easier.

myininaya (myininaya):

well you did more than I could I'm kinda like drawing a blank

geerky42 (geerky42):

just throw some thoughts. That's all.

OpenStudy (xapproachesinfinity):

hmm i see, sorry had some lag here lol

OpenStudy (xapproachesinfinity):

that one is wrong!

geerky42 (geerky42):

I know...

geerky42 (geerky42):

I somehow see \(5^n\), you know?

geerky42 (geerky42):

But you have any idea on \[\sum_{n=1}^{\infty}\frac{1}{(2n-1)^5}\] part?

OpenStudy (xapproachesinfinity):

i did see it that way first time lol

OpenStudy (xapproachesinfinity):

that still not easy to deal with lol

OpenStudy (xapproachesinfinity):

that's not a telescoping series is it?

OpenStudy (xapproachesinfinity):

@ganeshie8 see if you can help :)

geerky42 (geerky42):

I doubt it.

geerky42 (geerky42):

About telescoping serie lol

OpenStudy (xapproachesinfinity):

hmm integral test would work with that one

ganeshie8 (ganeshie8):

\[\begin{align}\sum_{1}^{\infty}(-1)^{n+1}\frac{1}{n^5} &=\sum_{1}^{\infty}\frac{1}{(2n-1)^5} - \frac{1}{(2n)^5}\\~\\ &=\sum_{1}^{\infty}\frac{1}{(2n-1)^5}\color{blue}{+\frac{1}{(2n)^5}-\frac{1}{(2n)^5}} - \frac{1}{(2n)^5}\\~\\ &=\sum_{1}^{\infty}\frac{1}{n^5} - \frac{2}{(2n)^5}\\~\\ &=\sum_{1}^{\infty}\frac{1}{n^5} - \frac{1}{16n^5}\\~\\ &=\frac{15}{16}\sum_{1}^{\infty}\frac{1}{n^5}\\~\\ &=\frac{15}{16}\zeta(5) \end{align}\]

geerky42 (geerky42):

Smart move.